What is the present value of a $100,000 cash flow to be received at the end of each of the next 15 years from an account that earns an annual rate of 10%?

1,500,000

760,608

523,450

976,455
Answer #1

Annual Cash Flows P = $100000

Interest Rate = r = 10%

Number of periods = n = 15

This is case of ordinary Annuity (since the payments are made at end of each year)

Hence, NPV = P/(1+r) + P/(1+r)2 +.... P/(1+r)15

=> NPV = P [1 - (1+r)-n]/r = 100000 (1 - 1.10-15)/0.10 = $760607.95

Hence the correct option is (b) $760,608